OpenStudy (anonymous):

Anyone know how to find the arc length of the curve: y=(((x^(4))/(4))+(1/(8x^(2))) on the interval 1less than or equal to x less than or equal to 2

OpenStudy (amistre64):

ds = sqrt(dx^2 + dt^2) if I recall correctly

OpenStudy (amistre64):

since y is a function of "x" dx^2 = 1

OpenStudy (amistre64):

dt was supposed to be dy.....

OpenStudy (anonymous):

\[L=\int\limits_{a}^{b}\sqrt{1+(dy/dx)^2 dx}\]

OpenStudy (anonymous):

a=1 b=2 but what is (dy/dx)?

OpenStudy (amistre64):

looks right.... but I think it has a "ds" on the end, not that it matters much :) and make sure your "dx" is out of the radical

OpenStudy (amistre64):

dy/dx = the derivative of your function

OpenStudy (anonymous):

I have (dy/dx)= x^(3) - (1/(4x^(3)))

OpenStudy (amistre64):

the length thing is the square root of the sum off the derivatives of x and y; since y is a function of x; x'=1

OpenStudy (anonymous):

Is that right?

OpenStudy (amistre64):

id have to try to decipher the original equation you posted and maybe implicit it for y'...

OpenStudy (anonymous):

okay hold on let me write it again, sorry for it being sloppy hahah

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

\[y=(x^{4}/4) + 1/(8x^2)\]

OpenStudy (anonymous):

A lil neater :D

OpenStudy (amistre64):

neater yes; but make sure that this is the correct equation.... double check it :)

OpenStudy (anonymous):

Yes, lol. And I have to find the arc length of the curve from \[1\le x \le 2\]

OpenStudy (amistre64):

4x^3 16x y' = ------ (-) --------- 4 (8x^2)^2 this should be y' = dy/dx

OpenStudy (amistre64):

can reduce as needed :)

OpenStudy (anonymous):

y' = x^(3) - (1/4x^(3)) reduced?

OpenStudy (amistre64):

yep, that looks good :)

OpenStudy (anonymous):

haha awesome.

OpenStudy (anonymous):

now I have to square that. That is when I got : x^(6) - 1/2 + 1/(16x^(6))

OpenStudy (amistre64):

as long as you kept track of all the stuff, ill trust you on that :)

OpenStudy (anonymous):

let me double check real quick, I'm not sureee..

OpenStudy (amistre64):

can we get that into 1 term? get like denominators be easier?

OpenStudy (anonymous):

I think that might make everything a bit more confusing... lol

OpenStudy (anonymous):

hold on a sec :)

OpenStudy (amistre64):

4x^6 - 1 -------- ?? 4x^3

OpenStudy (anonymous):

ummmm... I'm lost. Sorry! What did you do?

OpenStudy (anonymous):

nvm, I see what you did haha

OpenStudy (anonymous):

hm, maybe.. let me see real quick

OpenStudy (amistre64):

y' = x^(3) - (1/4x^(3)) get like denominators.... 4x^3 is common

OpenStudy (anonymous):

ahhhh, smart move, that helped. let me make that change real fast

OpenStudy (amistre64):

16x^12 -8x^6 +1 ------------------ = y'^2 16x^6

OpenStudy (anonymous):

Yeah! now I have to put that back into the equation for arc length..

OpenStudy (amistre64):

yep...

OpenStudy (amistre64):

reduce it if you want :)

OpenStudy (anonymous):

\[\int\limits_{1}^{2} x^6 + 1/16x^6 + 1/2 dx\]

OpenStudy (amistre64):

(S) sqrt[1 + x^6 + (1/16x^6) - (1/2)] dx

OpenStudy (anonymous):

oh thats not too bad

OpenStudy (amistre64):

dont forget its the "square root" of 1 + y'^2 ....

OpenStudy (anonymous):

ohhhhh, this is where I got caught up...

OpenStudy (anonymous):

yeah, thats where I didn't know what to do next. Should I take the square root of each term separately?

OpenStudy (amistre64):

cant do that; they aint multiplied together, it just a one lump sum under that radical and you cant break that up....

OpenStudy (anonymous):

x^(3) + 1/4x^(3) + 1/sqrt2 ?

OpenStudy (anonymous):

oh you cant do that :( darnnnn

OpenStudy (anonymous):

are you sure?

OpenStudy (amistre64):

watch: sqrt(4*3) = sqrt(4) * sqrt(3) = 2sqrt(3) BUT!! sqrt(4+3) NOT equal sqrt(4) + sqrt(3)

OpenStudy (anonymous):

ohhhhh, you're right. soooo... what do I do?

OpenStudy (anonymous):

:(

OpenStudy (amistre64):

I dont really know..... thats were I usually have to open up the book and re-read what they have in there for that section :)

OpenStudy (amistre64):

it looks like you might be able to do some sort of trig substitution, but i just aint confident about my abilities in that yet...

OpenStudy (anonymous):

ahhh, is there a quadratic function we can make out of those three terms?

OpenStudy (anonymous):

oh yeahhh, trig substitution... ugh

OpenStudy (amistre64):

sqrt(1+tan^2) = something or another.... but I always mess it up :)

OpenStudy (anonymous):

lol I'll check...

OpenStudy (anonymous):

thanks for helping me on such an exhausting problem! :D

OpenStudy (amistre64):

lol ..... I hope I learned something :)

OpenStudy (amistre64):

if y' = tan sqrt(1+tan^2) = sqrt(sec^2) = sec... (S) sec du or something...

OpenStudy (anonymous):

hahaha, ummm... I dont know if its a trig substitution. But if you wanted to move to another problem or anything thats fine. I dont want to bother you all night with just this problem

OpenStudy (anonymous):

totally your choice :)

OpenStudy (amistre64):

well, library is closing soon, which takes my internet with it :) So I will be heading out anywhoos :) It was fun... Ciao

OpenStudy (anonymous):

Thanks again :)